OpenStudy (anonymous):

Use the substitution method to solve the linear equation to find the solution. 7a= 5-b, 7a= 2-b. Will it be one solution,infinitely or on solution?

OpenStudy (loser66):

do you have any idea?

OpenStudy (pottersheep):

7a= 5-b, Can be rewritten as b= **5 - 7a** so if I sub that into 7a= 2-b [i have a value for b now] 7a = 2 - **[ 5 - 7a ]** 7a = 2 - [ 5 - 7a ] 7a = 2 - 5 + 7a move 7a to other side uh oh that makes 0 = 3 not possible. so no solution

OpenStudy (loser66):

congrat!

OpenStudy (loser66):

look at this, 7a = 5-b also, 7a = 2-b then --------------- 5-b= 2-b plus b to both sides 5-b+b= 2-b+b ----> 5=2 ???!!!! impossible ---> no solution